A number is five times another number. When the numbers are added, their sum is 24. What are the two numbers?
Answer by mananth(16949)

let one number be x
other is 5 times the number = 5x
...
add up
5x+x = 6x
6x = 24
/6
6x/6 = 24/6
x= 4 one number
other number = 24-4 = 20
..
m.ananth@hotmail.ca